if the length ia 4 yds longer double the witdth of a playing field perimetre is 266 yds what is the dimensions
Answer by checkley77(12569)

L=W+4
2L+2W=266
2(W+4)+2W=266
2W+8+2W=266
4W=266-8
4W=258
W=258/4
W=64.5 YDS IS THE WIDTH.
L=64.5+4
L=68.5 YDS IS THE LENGTH
PROOF
2*68.5+2*64.5=266
137+129=266
266=266